The average hospitality girls working in a certain area is 19 years old. A civic oriented group made a recent survey on the age of the hospitality girls working in the same locality. A random sample of 20 respondents showed an average of 15.4 years old with a standard deviation of 2.14 years old. Using a 0.05 level of significance, has the hospitality girls gone lower?

1
Expert's answer
2021-12-21T16:54:35-0500

The following null and alternative hypotheses need to be tested:

"H_0:\mu\geq 19"

"H_1:\mu<19"

This corresponds to a left-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=20-1=19" degrees of freedom, and the critical value for a left-tailed test is "t_c = -1.729133."

The rejection region for this left-tailed test is "R = \{t: t < -1.729133\}."

The t-statistic is computed as follows:


"t=\dfrac{X-\mu}{s/\sqrt{n}}=\dfrac{157.4-19}{2.14/\sqrt{20}}\approx-7.52322"

Since it is observed that "t = -7.52322<-1.729133=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach: The p-value for left-tailed, "df=19" degrees of fredom, "t=-7.52322" is "p \approx 0," and since "p=0<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"  is less than "19," at the "\alpha = 0.05" significance level.